Glycerol is a colorless, odorless, viscous liquid with a sweet taste and non-toxic in low concentration . It is obtain as a by product of manufactured soap. Its percent composition is 39.12% C, 8.75% hydrogen, and 52.12% oxygen. The molecular mass is 92.11 g/mol. What is the molecular formula for glycerol


1
Expert's answer
2021-05-24T07:19:45-0400

<div data-test="navigation_logo" class="sg-logo HeaderController__logo--27DEB" style="position: relative; overflow: hidden; text-overflow: ellipsis; line-height: 1.5rem; min-height: 1.5rem; margin: 0px; height: 32px; width: 72px; transition: opacity 0.25s ease 0s;"></div>


2

For studentsFor parentsFor teachersHonor codeBrainly Tutor

marywoozy

  • 2 weeks ago

ChemistryHigh School


answered

Glycerol has a molar mass of 92.09g/mol. Its percent composition is: 39.12% C, 8.75% H,and 51.12% O. What is the molecular formula for glycerol?



1

SEE ANSWER


ADD ANSWER

+5 PTS



<div class="sg-avatar__image sg-avatar__image--icon" style="display: flex; position: relative; height: 32px; width: 32px; background-color: rgb(104, 123, 140);"><div class="sg-icon sg-icon--gray-light sg-icon--x32 sg-avatar__icon" style="fill: rgb(195, 209, 221); max-width: 32px; max-height: 32px; height: 32px; width: 32px; flex-shrink: 0;"><svg class="sg-icon__svg"><use xlink:href="#icon-profile"></use></svg></div></div>

marywoozy is waiting for your help.

Add your answer and earn points.

Answer

5.0/5


2




Eduard22sly

  • Genius
  • 5.4K answers
  • 15.7M people helped

Answer:

Molecular formula => C₃H₈O₃

Explanation:

From the question given above, the following data were obtained:

Carbon (C) = 39.12%

Hydrogen (H) = 8.75%

Oxygen (O) = 51.12%

Molar mass of compound = 92.09 g/mol

Molecular formula =?

Next, we shall determine the empirical formula of the compound. This can be obtained as follow:

C = 39.12%

H = 8.75%

O = 51.12%

Divide by their molar mass

C = 39.12 / 12 = 3.26

H = 8.75 / 1 = 8.75

O = 51.12 / 16 = 3.195

Divide by the smallest

C = 3.26 / 3.195 = 1

H = 8.75 / 3.195 = 2.7

O = 3.195 / 3.195 = 1

Thus, the empirical formula is CH₂.₇O

Finally, we shall determine the molecular formula of the compound. This can be obtained as follow:

Empirical formula = CH₂.₇O

Molar mass of compound = 92.09 g/mol

Molecular formula =?

Molecular formula = Empirical formula × n

Molecular formula = [CH₂.₇O]ₙ

92.09 = [12 + (2.7×1) + 16] × n

92.09 = 30.7n

Divide both side by 30.7

n = 92.09 / 30.7

n = 3

Molecular formula = [CH₂.₇O]ₙ

Molecular formula = [CH₂.₇O]₃

Molecular formula = C₃H₈O₃